Ksp of Mg(OH)2 IS 4*10^-12 .the no. of moles of Mg2+ ions in one litre of its saturated solution in 0.1 M NaOH is??

Ksp of Mg(OH)2  IS 4*10^-12 .the no. of moles of Mg2+ ions in one litre of its saturated solution in 0.1  M NaOH  is??

Grade:12th pass

2 Answers

Arun
4 years ago
Ksp = [Mg2+] [OH¯]2
Let the concentration of Mg2+ ions to form saturated solution be s mol/litre
4 x 10-12 = s x (0.1)2 
s = 4 x 10-10 mol/litre
Volume of solution = 1 litre 
Number of moles of Mg2+ = Molarity x Volume = ​4 x 10-10 x 1 = ​4 x 10-10 mol
 
Vikas TU
4 years ago
Dear student 
Ksp refers to the solubility product constant. It refers to the equilibrium constant of a solid solid substance dissolving in aqueous solution. The more soluble a substance is the higher the Ksp of that given solid. 
so with that background knowledge, we can now proceed to the calculations;
first write down the equation; Mg(OH)2  ⇒ Mg2+ + 2OH-
                            Ksp= [Mg2+] [OH-]²
                       4.0x10^-12= S x (2S)²     where S is solubility
              therefore 4.0 x 10-12= 4S³
                        4.0 x 10-12= 0.1(4S²)
               S= 3.16 x 10^-6
    hence concentration of NaOH= 2 x (3.16 x 10^-6)
                                                  = 6.32 x 10^-6